OpenStudy (anonymous):

How would you graph P(x)=32-2x^4

OpenStudy (campbell_st):

take 2 out as a common factor P(x) = 2(16 - x^4) let P(x) = 0 and so solve 0 = 2(16 - x^4) this is an equation that can be reduced to a quadratic \[0 = 2(16 - (x^2)^2)\] so you are dealing with the difference of 2 squares. hope this helps

OpenStudy (anonymous):

ok, so i simplified it to 2(2-x)(2+x)(4+x^2) and i got the x-intercepts to be 2, -2, 2i and -2i. is this right and if so how would you graph imaginaries?

OpenStudy (campbell_st):

next you can differentiate to find any stationary points... and then find the 2nd derivative to find any possible points of inflexion and to also test the stationary points, check for maximums, minimums or horizontal points of inflexion

OpenStudy (anonymous):

i havent learned derivatives yet, im only halfway through algebra 2. is there any way to do it that my instructor would recognize as what we are learning

OpenStudy (campbell_st):

you'll just graph the real zeros..

OpenStudy (anonymous):

ok so just x intercepts os 2 and -2 and y int of 32?

OpenStudy (campbell_st):

ok... here is a simple method, use a table of values  when you have the table plot the points on a number plane

OpenStudy (campbell_st):

this is a parabola shape but steeper, and concave down, Hope this helps

OpenStudy (anonymous):

thanks